How much heat is necessary to change 500g of ice at 263K to water at 293K
1
Expert's answer
2021-09-03T13:56:29-0400

"m=500 \;g \\

T_1= 263-273 = -10 \;°C \\

T_2 = 293 - 273 = 20 \; °C"

Specific heat capacity for Ice = 2.1 J/g°C

Latent heat of fusion of Ice = 335 J/g

Specific heat capacity for Water = 4.2 J/g°C

"Q_1 = 10\times 500 \times 2.1 = 10500 \;J" (from -10 °C to 0 °C)

"Q_2 = 20 \times 500 \times 4.2 = 42000 \;J" (from 0 °C to 20 °C)

"Q_3 = 500 \times 335 = 167500 \;J" (heat of melting)

Total heat "Q_T = 10500+42000+167500 = 220000\;J"

Answer: 220 kJ